OpenStudy (anonymous):

A particle in free-fall: a) takes 42 seconds to travel 490 meters at g=9.8 m/s2 B) takes 15.8 seconds to travel 490 meters at g= 9.8 m/s2 C) takes 4.5 seconds to travel 490 meters at g= 9.8 m/s2 D) takes 10 seconds to travel 490 meters at g= 9.8 m/s2

OpenStudy (shamim):

Given u=0 h=490m g=9.8m/s^2 t=?

OpenStudy (shamim):

Use h=ut+.5gt^2

OpenStudy (anonymous):

ok thanks

OpenStudy (shamim):

Can u tell me the right option?

OpenStudy (shamim):

Anyway welcome!!

OpenStudy (shamim):

No its not B

OpenStudy (anonymous):

Vf = Vi + at you know everything except t. solve for t.

OpenStudy (anonymous):

sorry, also use vf^2 - vi^2 = 2aX

OpenStudy (anonymous):

the answer is not B...its 9.9 sec so its D